OpenStudy (anonymous):

Supply the missing term or number to make a perfect square trinomial.y^2 + 8y + ______

OpenStudy (raden):

add by 1/2 cofficient of y then square it so, add by (1/2 * 8)^2 = 4^2 = 16

OpenStudy (anonymous):

Both are correct answers. But since I don't know alot about this at all, looking for the best way to get the answer needed. I'm totally lost on most of it. Sorry but thanks for the help.